# Adversarial Alignment: Ensuring Value Consistency in Large Language Models for Sensitive Domains

### v1

@@ Abstract
keywords:
Large Language Models; Adversarial Alignment; Value Consistency
journal: Data Intelligence
@@ 3.2 Adversarial Training
The prompt templates for the Attacker, Actor, and Critic are shown in A.
@@ 3.3 VC-LLM
Details of the training data and hyperparameters are provided in B.
@@ 4.1 Experimental Setups
We construct a bilingual value evaluation dataset (Chinese and English), covering five sensitive domains: Sovereignty (Sov), Human Rights (HR), Religion (Rel), Politics (Pol), and Ethnicity (Eth), further divided into 31 subcategories, with details in C.
@@ 4.3 Effectiveness of Adversarial Training
Table 4: Performance comparison of subdomains within Sovereignty in the Chinese evaluation dataset (average scores; full forms of abbreviations in the table are provided in C).
@@ Appendix A Prompt of The Attacker, Actor and Critic
Figure 1: Prompt of the Attacker, Critic and Actor. This data generation method may contain harmful content and is intended for research purposes only. Any other use is strictly prohibited.
@@ Appendix B Source and size of continued pre-training and fine-tuning corpus
Table 1: Source and size of continued pre-training corpus.
@@
Table 2: Source and size of instruction fine-tuning corpus.
@@
Table 3: Hyperparameter configurations used in the three training stages: continued pre-training, instruction fine-tuning, and adversarial training. All stages employ LoRA-based parameter-efficient adaptation.
@@ Appendix C Subdomains of Value Alignment Evaluation Dataset
Table 4: Subdomains of value alignment evaluation dataset. Warning: The dataset is for assessing the model’s value consistency only. The authors do not endorse any viewpoints in the examples that conflict with their values. The content is for research purposes only and must not be used otherwise.

### v2

@@ Abstract
Keywords Large Language Models; Adversarial Alignment; Value Consistency
@@ 3.2 Adversarial Training
The prompt templates for the Attacker, Actor, and Critic are shown in AppendixA.
@@ 3.3 VC-LLM
Details of the training data and hyperparameters are provided in AppendixB.
@@ 4.1 Experimental Setups
We construct a bilingual value evaluation dataset (Chinese and English), covering five sensitive domains: Sovereignty (Sov), Human Rights (HR), Religion (Rel), Politics (Pol), and Ethnicity (Eth), further divided into 31 subcategories, with details in AppendixC.
@@ 4.3 Effectiveness of Adversarial Training
Table 4: Performance comparison of subdomains within Sovereignty in the Chinese evaluation dataset (average scores; full forms of abbreviations in the table are provided in AppendixC).
@@ Appendix A Prompt of The Attacker, Actor and Critic
Figure A.1: Prompt of the Attacker, Critic and Actor. This data generation method may contain harmful content and is intended for research purposes only. Any other use is strictly prohibited.
@@ Appendix B Source and size of continued pre-training and fine-tuning corpus
Table B.1: Source and size of continued pre-training corpus.
@@
Table B.2: Source and size of instruction fine-tuning corpus.
@@
Table B.3: Hyperparameter configurations used in the three training stages: continued pre-training, instruction fine-tuning, and adversarial training. All stages employ LoRA-based parameter-efficient adaptation.
@@ Appendix C Subdomains of Value Alignment Evaluation Dataset
Table C.1: Subdomains of value alignment evaluation dataset. Warning: The dataset is for assessing the model’s value consistency only. The authors do not endorse any viewpoints in the examples that conflict with their values. The content is for research purposes only and must not be used otherwise.